what is
1. 2(3x-5)+2(60)= 180
2. Find the area of a rectangluar tennis court if the perimeter is 480 feet. the length of the court is six feet more than twice the width.
i dont understand please help!!!

Answer by tran3209(100)

The first Problem
2(3x-5) + 2(60)= 180
6x-10 + 120=180
6x+110= 180
6x=70
x=70/6
x = approx. 11.67
Check.
2(3*11.67-5) + 2(60) = 180
2(30.01) + 120 +180

60.02 + 120 = 180
180.02 = 180
180=180
_______________________________________________________________________________
Problem 2
The perimeter of the whole court is 480 ft.
Suppose "x" is the width of the court.
the length of the court would be
2x+6
To find perimeter of a rectangle the formula is 2l +2w
2x + 2(2x+6) = 480
2x + 4x + 12 =480
6x + 12 = 480
6x = 468
x= 468/6
x=78 ft
The width is 78 ft.
The length is 90 ft.
Check.
2(78) + 2(90) = 480ft^2